# EDA-Driven Preprocessing for SAT Solving

### v1

@@ Abstract
Effective formulation of problems into Conjunctive Normal Form (CNF) is critical in modern Boolean Satisfiability (SAT) solving for optimizing solver performance. Addressing the limitations of existing methods, our Electronic Design Automation (EDA)-driven preprocessing framework introduces a novel methodology for preparing SAT instances, leveraging both circuit and CNF formats for enhanced flexibility and efficiency. Central to our approach is the integration of a new logic synthesis technique, guided by a reinforcement learning agent, and a novel cost-customized LUT mapping strategy, enabling efficient handling of diverse SAT challenges. By transforming the SAT competition benchmarks into circuit instances, our framework demonstrates substantial performance improvements, as evidenced by a $52.42\%$ reduction on average compared to solving directly. Moreover, our framework achieves a remarkable $96.14\%$ runtime reduction on average for a set of logic equivalence checking problems that exhibit inherent circuit structures. These results highlight the effectiveness and versatility of our approach in handling both CNF and circuit instances. The code is available at https://github.com/cure-lab/EDA4SAT.
keywords:
SAT Preprocessing, Logic Synthesis, Circuit Learning
1 Introduction
In the realm of computational problem-solving, the Boolean Satisfiability (SAT) problem holds a pivotal position, which involves determining the truth of a Boolean formula under variable assignments. Identified as the first NP-complete problem [10], SAT’s complexity and versatility have captivated researchers for decades. SAT solvers [36, 16], central in addressing these problems, predominantly rely on Conjunctive Normal Form (CNF). CNF represents problems as an intersection of simpler clauses, a format that both challenges and drives innovation in SAT solving.
The dominance of CNF in SAT solving necessitates the conversion of diverse SAT problems into this format, be they graph-based, circuit-based, or pure Boolean formulas. Research spanning decades highlights the profound impact of CNF representation on problem-solving efficiency [41, 38]. A variety of preprocessing methods have emerged to refine CNF construction and transformation, optimizing the initial stages of the SAT solving process [12, 3, 39, 9]. Despite these developments, current SAT solving approaches continue to encounter significant challenges, underscoring the need for further innovation.
A major challenge within the existing SAT solving framework is the absence of a universally efficient formulation mechanism. In practical scenarios, SAT problems are often directly transformed into SAT instances, bypassing optimized constructions. This unrefined mapping often results in significantly larger problem sizes [3]. While various preprocessing techniques have been developed to compress the CNF formulae [30, 14, 37, 15], their effectiveness is not guaranteed. For instance, even seemingly small verification tasks can require hours to resolve [27], suggesting that simple CNF reduction may not suffice for efficient problem solving [9].
Moreover, several SAT preprocessing methods have been tailored for specific applications, such as microprocessor verification [39] and scheduling problems [12]. Although these methods perform well within their respective domains, their lack of adaptability to other problem types reveals a significant gap in the current approach to SAT preprocessing, highlighting the need for more universally applicable solutions.
To address these challenges, we introduce an Electronic Design Automation (EDA)-driven preprocessing framework, a novel integration into the standard SAT solving pipeline. This framework begins by converting CNF into a circuit format and then strategically reformulates the circuit representation. The process culminates in reverting the circuit back to an optimized CNF format, tailored for more efficient solving. This method’s flexibility in transitioning between circuit and CNF formats enables the application of advanced circuit transformation techniques across diverse SAT problems.
In our framework, we leverage EDA solutions to harness the detailed structural information of circuits for SAT problem reformulation. Central to this approach is the use of logic synthesis, augmented by a reinforcement learning (RL) agent dedicated to identifying optimal logic synthesis strategies. In contrast to traditional methods [15] focusing primarily on size reduction, our RL agent is specifically trained to minimize overall SAT solving complexity. The RL agent utilizes DeepGate [34], a sophisticated gate-level representation technique, for capturing rich information from initial problem instances. Additionally, we integrate a cost-customized look-up table (LUT) mapping strategy. Again, our approach maps the circuit into a format that is inherently easier to solve, diverging from conventional mappings aimed at compact design.
By integrating this EDA-driven preprocessing framework with the baseline SAT solver Kissat [16], we have achieved a remarkable average solving time reduction of $96.14\%$ for a set of circuit equivalence checking problems, compared to applying the baseline solver directly. Furthermore, we convert the CNF instances into circuits and then process circuits by our framework. Our framework demonstrates impressive versatility in effectively accelerating the solving process for a variety of problems in the SAT competition benchmarks with an average reduction in runtime of $52.42\%$.
We summarize the contributions of this work as follows:
  • Introduction of an innovative EDA-driven preprocessing framework that seamlessly integrates into the SAT solving pipeline, efficiently reformulating SAT problems by converting standard CNF formulas into circuits and optimizing them for easier solving. Additionally, the proposed framework is not mutually exclusive with the existing CNF-based preprocessing strategy.
  • Implementation of logic synthesis techniques guided by an RL agent, designed to explore optimal logic synthesis strategies and directly minimize the solving complexity.
  • Integration of a cost-customized LUT mapping approach that not only conceals redundant logic during solving, but also significantly enhances the reduction of solving complexity.
2 Related Work
@@
2.1 SAT Preprocessing
Conjunctive Normal Form (CNF) has become a standard format in modern SAT solvers, capable of representing arbitrary Boolean formulas. It is structured as a conjunction of clauses, denoted by $\phi=(C_{1}\land C_{2}\land\ldots)$, with each clause being a disjunction of variables or their negations, e.g., $C_{i}=(x_{1}\lor\neg x_{2}\lor\ldots)$. Over recent decades, numerous CNF-based heuristics have been developed to enhance SAT solving efficiency, including advanced clause management strategies [1] and innovative branching heuristics [29, 33].
While some studies [39, 12] have successfully modeled practical problems into CNFs, their specificity often limits generalizability to diverse problem sets. Commonly, real-world problems are converted directly into CNF, leading to redundant clauses and increased complexity. For example, a tournament scheduling problem with $n$ teams can escalate to $n^{6}$ clauses [3].
There are some works dedicated to CNF reformulation [14, 37], but the flattened structure and the lack of semantic information of CNF often require repetitively resolving clauses to establish logical relationships, which is time-intensive. Efforts like [30, 15] target recovering structures from CNF and minimizing the instances, but they cannot promise to reduce SAT solving time. In our work, we introduce a comprehensive preprocessing framework designed to directly reduce solving time.
2.2 Boolean Circuit Optimization
Boolean circuits, representable as directed acyclic graphs (DAGs) $\mathcal{G}(\mathcal{V},\mathcal{E})$ with logic gates and wires serving as nodes $\mathcal{V}$ and edges $\mathcal{E}$, respectively, are pivotal in computational logic. In EDA community, a wealth of efficient algorithms has been developed for circuit optimization. Logic synthesis is a prime example, involving DAG-based transformation techniques to modify circuit’s local structure for the optimization of performance, power, and area (PPA).
Figure 2: Example of logic rewrite and balance
We provide the examples of two logic synthesis methods that are supported by common EDA tools [7, 40]: rewrite and balance (see Figure 2). Logic rewrite (rewrite) is a circuit optimization technique that aims to minimize the overall circuit structure [6, 26]. The rewrite tools pre-compute and store all optimal circuits within $4$ inputs. When applied to an arbitrary circuit, the tools iteratively identify sub-graph and subsequently replace it with functionally-equivalent but smaller sub-graph. Balance (balance[11] creates an equivalent circuit with the minimum possible delay, i.e., the minimum number of logic levels. As suggested by its name, this method achieves optimization by balancing the fan-in regions of multi-input gates within the circuit.
@@ 2.2 Boolean Circuit Optimization
In this work, we introduce a novel approach where logic synthesis techniques are employed to simplify the solving complexity of circuit instances, rather than conventional goal of optimizing area or delay. Our proposed framework integrates with open-source EDA tool ABC [7] to perform logic synthesis and utilize customized LUT mapping to generate simplified circuit instances composed of LUTs.
2.3 Circuit Representation Learning
Training a pre-trained model to learn general representations and subsequently fine-tuning it for specific tasks has become a prominent trend across various domains. Recently, there has been a growing focus on developing pre-trained models capable of deriving circuit general representations. DeepGate family [21, 34] is a series of circuit representation learning models. The first version DeepGate [21] proposes to introduce strong inductive bias into circuit data through logic synthesis and is supervised by the statistic random simulation results, which effectively reflect structural and functional features of circuits. DeepGate2 [34] further advances circuit representation learning by disentangling the structural and functional representations. Moreover, DeepGate2 showcases significantly faster inference runtimes compared to the predecessor while even learning better circuit representations. The gate-level representation vectors derived from the DeepGate model encapsulate the general circuit information, which proves instrumental in facilitating various EDA applications, such as testability analysis [32] and power estimation [20]. In this work, we tailor existing EDA tools for specific requirements, utilizing circuit knowledge learning capabilities of pre-trained DeepGate2 to inform our circuit optimization decisions.
3 Methodology
3.1 Overview of the Proposed Framework
Figure 1 shows the overview of proposed EDA-Driven SAT Preprocessing. Firstly, we convert the initial CNFs into And-Inverter Graph (AIGs) [5], which represent any logic behaviours only with three gate types: AND Gate, NOT Gate and Primary Input (PI). Then, we apply a predetermined sequence of AIG transformation operations to unify the distribution of input circuit [22].
@@ 3.1 Overview of the Proposed Framework
Secondly, we utilize logic synthesis technique to reformulate SAT instance. Since the logic synthesis process can be considered as a sequential decision-making task, we train an agent to minimize the solving complexity by reinforcement learning (RL). In the RL framework, the state comprises embeddings of the initial AIG netlist and features of intermediate circuits. Additionally, the RL agent operates within a discrete action space, where each action corresponds to executing a synthesis operation. We approximate the complexity of SAT solving based on the number of variable branching times and define the reduction in branching times as the reward for RL agent.
@@
Thirdly, we employ the LUT mapping technique to hide the internal logic of circuits. To further decrease the solving time, we design a cost-customized mapping operation with the objective of minimizing the branching times during SAT solving, rather than solely focusing on reducing the problem size. After the LUT mapping, we obtain the LUT netlist with a relatively small number of cells. To ensure compatibility with modern SAT solvers, we transform the LUT netlist back into a simplified CNF.
3.2 Logic Synthesis Recipe Exploration
@@
We formulate logic synthesis as a Markov Decision Process (MDP) process, with the optimization objective of reducing the solving time for a given input instance. Consequently, such sequence decision-making process entails iteratively selecting a logic synthesis operation as an action according to the current circuit state.
3.2.1 Reinforcement Learning Formulation
@@
To explore the optimal logic synthesis sequence (i.e. logic synthesis recipe), we employ Deep Q-learning algorithm in this section to train an reinforcement learning (RL) agent. The RL agent is trained to make informed decisions to minimize solving time by learning from its interactions with the circuit environment.
@@
As shown in Figure 3, at each step $t$ (where $t=0,1,...,T-1$), the features of the current netlist $\mathcal{G}^{t}$ are extracted to form the state $s^{t}$. Subsequently, the Q-learning agent utilizes $s^{t}$ as input and selects an action $a^{t}$. Then, the environment including the logic synthesis tool performs the synthesis operation based on $a^{t}$ and transforms the netlist $\mathcal{G}^{t}$ into a new netlist $\mathcal{G}^{t+1}$. At the end of this step, the environment provides RL agent with a reward $r^{t}$. The above process is repeated for subsequent steps until $t=T-1$. During the training process, the objective of RL agent is to maximize the cumulative sum of the reward values. As a result, the trained RL agent is able to select logic synthesis operations step by step to achieve the optimal solving time reduction.
@@ 3.2.2 State
  • The ratio between the area of netlist $\mathcal{G}^{t}$ and netlist $\mathcal{G}^{0}$.
  • The ratio between the depth of netlist $\mathcal{G}^{t}$ and netlist $\mathcal{G}^{0}$.
  • The ratio between the wire counts of netlist $\mathcal{G}^{t}$ and netlist $\mathcal{G}^{0}$.
  • The proportion of AND gates in the total gates of netlist $\mathcal{G}^{t}$.
  • The proportion of NOT gates in the total gates of netlist $\mathcal{G}^{t}$.
  • The average balance ratio of netlist $\mathcal{G}^{t}$. We formally define the average balance ratio $br$ in Eq. (1), where $\mathcal{P}_{i1}$, $\mathcal{P}_{i2}$ are predecessors of two fanin AND gate $i$, and $d$ denotes the depth.
@@
3.2.3 Action
The action space $\mathcal{A}$ of agent is discrete and encompasses logic synthesis operations. In this paper, we set the available operations as follows: rewrite [26], refactor [8], balance [11], resub [7] and end (marking the end of the logic synthesis process). The selection of these operations does not imply that the proposed framework lacks support for additional synthesis operations beyond this set. Instead, these operations are chosen due to their widespread usage and commonality. If there is a need to include additional optional operations, it can be accomplished by expanding the action space within the framework.
3.2.4 State Transition
The state transition function is implemented by the logic synthesis tool. We call the tool with a specific logic synthesis operation determined by RL agent to transform current netlist $\mathcal{G}^{t}$ to another functional-equivalent but simplified netlist. More formally, we note the given netlist as $\mathcal{G}^{t}$ and RL action as $a^{t}$ at step $t$. The state transition function is denoted as: $\mathcal{G}^{t+1}=\mathcal{F}(\mathcal{G}^{t},a^{t})$, where $\mathcal{G}^{t+1}$ is the updated netlist by tool. The updated netlist, $\mathcal{G}^{t+1}$, is then considered as the input for the subsequent decision step at time $t+1$.
3.2.5 Reward
@@
Prior to solving the SAT instance in circuit format, we employ the LUT mapping technique to convert the AIG netlist into an LUT netlist. Subsequently, the transformed netlist is converted into Conjunctive Normal Form (CNF), which will be elaborated in the following section. Additionally, we choose not to minimize the solver runtime based on two primary reasons. Firstly, determining the solver runtime for hard SAT instances is significantly time-consuming. Secondly, the solver runtime is challenging to calculate accurately due to the fluctuating CPU status in the case of easy SAT instances. We utilize variable branching times to approximate solving time. This approach allows us to include easy cases with short solver runtime in the training dataset, thereby enhancing the efficiency of training the RL agent.
3.2.6 Policy
In the Q-learning framework [28], the policy for maximizing the total reward entails selecting the action with the highest expected value. We implement the action-value function $Q$ using multilayer perceptron (MLP) to estimate the sum of future rewards after taking a certain action under current state. The policy $\pi(s^{t})$ and action-value function $Q_{\theta}(s^{t},a)$ are shown in Eq. (4), where the $\theta$ parametrizes the function $Q$.

### v2

# Logic Optimization Meets SAT: A Novel Framework for Circuit-SAT Solving

@@ Abstract
The Circuit Satisfiability (CSAT) problem, a variant of the Boolean Satisfiability (SAT) problem, plays a critical role in integrated circuit design and verification. However, existing SAT solvers, optimized for Conjunctive Normal Form (CNF), often struggle with the intrinsic complexity of circuit structures when directly applied to CSAT instances. To address this challenge, we propose a novel preprocessing framework that leverages advanced logic synthesis techniques and a reinforcement learning (RL) agent to optimize CSAT problem instances. The framework introduces a cost-customized Look-Up Table (LUT) mapping strategy that prioritizes solving efficiency, effectively transforming circuits into simplified forms tailored for SAT solvers. Our method achieves significant runtime reductions across diverse industrial-scale CSAT benchmarks, seamlessly integrating with state-of-the-art SAT solvers. Extensive experimental evaluations demonstrate up to 63% reduction in solving time compared to conventional approaches, highlighting the potential of EDA-driven innovations to advance SAT-solving capabilities.
I Introduction
The Circuit Satisfiability (CSAT) problem lies at the heart of many critical tasks in Electronic Design Automation (EDA), such as logic equivalence checking (LEC), automatic test pattern generation (ATPG), and logic synthesis (LS). It involves determining whether a Boolean circuit has an input assignment that satisfies its primary output. As a specialized variant of the Boolean Satisfiability (SAT) problem, CSAT is known to be NP-complete [1], making it both theoretically challenging and practically important in integrated circuit (IC) design and verification.
Modern SAT solvers, such as MiniSAT, Kissat, and CaDiCaL [2, 3], have been highly effective in solving problems represented in Conjunctive Normal Form (CNF). However, when applied to CSAT problems, these solvers face inefficiencies stemming from the required transformation of circuits into CNF format. This transformation often disrupts the natural structure of circuits, leading to representations that are less solver-friendly and more complex to process [4, 5]. As a result, solving CSAT instances may become unnecessarily time-consuming, particularly for large-scale, industrial circuits.
While preprocessing techniques for CNF have been extensively explored [4, 5, 6, 7], similar advancements tailored specifically for circuit-based representations remain underdeveloped. Circuit-specific methods often focus on reducing circuit size or area [8], which, while useful for design optimization, do not directly address the complexity of SAT solving. This gap highlights the need for a new approach that bridges the structural strengths of circuit representations with the solving power of modern SAT solvers.
To address these limitations, we propose a novel preprocessing framework tailored to CSAT problems. Our approach integrates EDA techniques with reinforcement learning (RL) to transform CSAT instances into SAT-friendly formats. By leveraging logic synthesis tools and a cost-customized Look-Up Table (LUT) mapping strategy, our framework achieves significant reductions in solving complexity. Unlike traditional preprocessing methods, our approach explicitly prioritizes SAT-solving efficiency rather than widely-used optimization metrics like area or delay. Specifically, the main contributions of our work include:
  • EDA-Driven Preprocessing Framework: We introduce a comprehensive preprocessing pipeline that combines logic synthesis techniques with reinforcement learning to optimize CSAT instances for efficient SAT solving. This integration bridges the gap between raw circuit structures and modern SAT solvers.
  • RL-Guided Logic Synthesis for SAT Solving: We formulate logic synthesis as a Markov Decision Process (MDP) and train an RL agent to identify synthesis sequences that minimize solving complexity. The agent leverages circuit features inspired by [9, 10] and structural circuit embeddings from DeepGate2 [11].
  • Cost-Customized LUT Mapping for SAT Solving: We develop a novel LUT mapping strategy that incorporates a customized cost metric reflecting branching complexity during SAT solving. This approach diverges from conventional LUT mapping methods [12, 13] by focusing on solving efficiency.
Using industrial-scale benchmarks derived from LEC and ATPG tasks, we demonstrate significant runtime reductions up to 63% compared to baseline approaches. Our method integrates seamlessly with state-of-the-art SAT solvers, such as Kissat and CaDiCaL.
II Related Work
@@
II-A SAT and Circuit SAT Problems
SAT problem holds a pivotal position in hardware design, software verification, planning and scheduling, which seeks to identify if there exists at least one assignment that makes a given Boolean formula True. The modern SAT solvers, central in addressing SAT problems, predominantly rely on the Conflict-Driven Clause Learning (CDCL) algorithm [14]. Within the backbone CDCL algorithm, CNF has become a standard format in SAT solving, capable of representing arbitrary Boolean formulas. It is structured as a conjunction of clauses, denoted by $\phi=(C_{1}\land C_{2}\land\ldots)$, with each clause being a disjunction of variables or their negations, e.g., $C_{i}=(x_{1}\lor\neg x_{2}\lor\ldots)$. Over recent decades, numerous CNF-based heuristics have been developed to enhance SAT solving efficiency, including advanced clause management strategies [15] and innovative branching heuristics [16, 17].
Typically, practical CNF-modeled problems often result in redundant clauses and increased complexity. For example, a tournament scheduling problem with $n$ teams can escalate to $n^{6}$ clauses [18]. Consequently, problem reformulation, as known as preprocessing, is essential before solving. Efforts like [5, 6, 4] target minimizing the CNF instances by resolution and graph transformation, resulting in faster SAT solving.
One of the most representative SAT variants in the realm of EDA is the Circuit SAT (CSAT) problem. In contrast to the general SAT problem, the CSAT problem determines the satisfiability of a Boolean circuit. Over the past decades, circuit-based SAT solvers have been developed to tackle the CSAT problems [19, 20, 21]. Despite these solvers conducting logic reasoning in natural formats, they still exhibit lower performance levels compared to CNF-based SAT solvers due to the inefficient circuit-based preprocessing and heuristics. For example, the previous circuit-based preprocessing approaches, such as  [8], cannot promise to reduce solving time as they only scale down circuits. In this work, we introduce an efficient and effective circuit-based preprocessing framework to directly mitigate solving complexity by integrating EDA tools and reinforcement learning guidance.
II-B Boolean Circuit Optimization
Boolean circuits, representable as directed acyclic graphs (DAGs) $\mathcal{G}(\mathcal{V},\mathcal{E})$ with logic gates and wires serving as nodes $\mathcal{V}$ and edges $\mathcal{E}$, respectively, are pivotal in computational logic. In the EDA community, a wealth of efficient algorithms has been developed for circuit optimization. Logic Synthesis (LS) is a prime example, involving DAG-based transformation techniques to modify circuit’s local structure for the optimization of performance, power, and area (PPA).
@@ II-B Boolean Circuit Optimization
The advent of circuit representation learning [23] marks a progressive trend in the field, focusing on deriving neural representations of circuits to facilitate downstream tasks. The DeepGate family [24, 11, 25], for instance, integrates structural and functional aspects of circuits and has shown impressive results in testability analysis [26] and SAT solving [21]. In our research, we tailor existing EDA tools for specific requirements, utilizing DeepGate2’s [11] circuit knowledge learning capabilities to inform our circuit optimization decisions.
III Methodology
III-A Overview of the Proposed Framework
Fig. 1 shows the overview of the proposed circuit-based preprocessing framework. Given a CSAT problem represented in And-Inverter Graphs (AIGs), which only contain three gate types (primary input, AND and NOT) by [27], the baseline pipeline converts it into CNF.
@@ III-A Overview of the Proposed Framework
In our proposed framework, we first apply a predetermined sequence of AIG transformation operations to unify the distribution of input circuits [21]. Then, we utilize logic synthesis technique to reformulate SAT instance. Since the logic synthesis process can be considered as a sequential decision-making task, we train an agent to minimize the solving complexity by reinforcement learning (RL). In the RL framework, the state comprises embeddings of the initial AIG netlist and features of intermediate circuits. Additionally, the RL agent operates within a discrete action space, where each action corresponds to executing a synthesis operation. We approximate the complexity of SAT solving based on the number of variable branching times and define the reduction in branching times as the reward for RL agent.
@@
Finally, we employ the LUT mapping technique to hide the internal logic of circuits. To further decrease the solving time, we design a cost-customized mapping operation with the objective of minimizing the branching times during SAT solving, rather than solely focusing on reducing the problem size. After the LUT mapping, we obtain the LUT netlist with a relatively small number of cells. To ensure compatibility with modern SAT solvers, we transform the LUT netlist back into a simplified CNF.
III-B Logic Synthesis Recipe Exploration
@@
We formulate LS as a Markov Decision Process (MDP) process, with the optimization objective of reducing the solving time for a given input instance. Consequently, such sequential decision-making process entails iteratively selecting a LS operation as an action according to the current circuit state.
III-B1 Reinforcement Learning Formulation
@@
To explore the optimal LS sequence (i.e. LS recipe), we employ Deep Q-learning algorithm in this section to train a reinforcement learning (RL) agent. The RL agent is trained to make informed decisions to minimize solving time by learning from its interactions with the circuit environment.
@@
As shown in Fig. 2, at each step $t$ (where $t=0,1,...,T-1$), the features of the current netlist $\mathcal{G}^{t}$ are extracted to form the state $s^{t}$. Subsequently, the Q-learning agent utilizes $s^{t}$ as input and selects an action $a^{t}$. Then, the environment including the LS tool performs the synthesis operation based on $a^{t}$ and transforms the netlist $\mathcal{G}^{t}$ into a new netlist $\mathcal{G}^{t+1}$. At the end of this step, the environment provides RL agent with a reward $r^{t}$. The above process is repeated for subsequent steps until $t=T-1$. During the training process, the objective of RL agent is to maximize the cumulative sum of the reward values. As a result, the trained RL agent is able to select LS operations step by step to achieve the optimal solving time reduction.
@@ III-B2 State
  • Ratio between the area of netlist $\mathcal{G}^{t}$ and netlist $\mathcal{G}^{0}$.
  • Ratio between the depth of netlist $\mathcal{G}^{t}$ and netlist $\mathcal{G}^{0}$.
  • Ratio between the wire counts of netlist $\mathcal{G}^{t}$ and netlist $\mathcal{G}^{0}$.
  • Proportion of AND gates in the total gates of netlist $\mathcal{G}^{t}$.
  • Proportion of NOT gates in the total gates of netlist $\mathcal{G}^{t}$.
  • Average balance ratio of netlist $\mathcal{G}^{t}$. We formally define the average balance ratio $br$ in Eq. (1), where $\mathcal{P}_{i1}$, $\mathcal{P}_{i2}$ are predecessors of two fanin AND gate $i$, and $d$ denotes the depth.
@@
III-B3 Action
The action space $\mathcal{A}$ of agent is discrete and encompasses LS operations. In this paper, we set the available operations as follows: rewrite [28], refactor [29], balance [30], resub [31] and end (marking the end of the LS process). The selection of these operations does not imply that the proposed framework lacks support for additional synthesis operations beyond this set. Instead, these operations are chosen due to their widespread usage and commonality. If there is a need to include additional optional operations, it can be accomplished by expanding the action space within the framework.
III-B4 State Transition
The state transition function is implemented by the LS tool. We call the tool with a specific LS operation determined by RL agent to transform the current netlist $\mathcal{G}^{t}$ to another functional-equivalent but simplified netlist. More formally, we note the given netlist as $\mathcal{G}^{t}$ and RL action as $a^{t}$ at step $t$. The state transition function is denoted as: $\mathcal{G}^{t+1}=\mathcal{F}(\mathcal{G}^{t},a^{t})$, where $\mathcal{G}^{t+1}$ is the updated netlist by tool. The updated netlist, $\mathcal{G}^{t+1}$, is then considered as the input for the subsequent decision step at time $t+1$.
III-B5 Reward
@@
Prior to solving the SAT instance in circuit format, we employ the LUT mapping to convert the AIG netlist into an LUT netlist. Subsequently, the transformed netlist is converted into CNF, which will be elaborated in the following section. Additionally, we choose not to minimize the solver runtime based on two primary reasons. Firstly, determining the solver runtime for hard SAT instances is time-consuming. Secondly, the solver runtime is challenging to calculate accurately due to the fluctuating CPU status in the case of easy SAT instances. We utilize variable branching times to approximate solving time. This approach allows us to include easy cases with short solver runtime in the training dataset, thereby enhancing the efficiency of training the RL agent.
III-B6 Policy
In the Q-learning framework [32], the policy for maximizing the total reward entails selecting the action with the highest expected value. The action value function $Q$ uses the multilayer perceptron (MLP) to estimate the sum of future rewards after taking a certain action in the current state. The policy $\pi(s^{t})$ and action-value function $Q_{\theta}(s^{t},a)$ are shown in Eq. (4), where the $\theta$ parametrizes the function $Q$.